A piece of silver wire has a resistance of 1.0Ω. What will be the resistance of a manganin wire; 

one-third the length and one-third the diameter if the resistivity of manganin is 30 times that of silver.


1
Expert's answer
2021-09-29T18:37:12-0400

For aluminium wire

Resistance=R

Length=L

Resistivity="\rho"

Diameter=D

Area (A)="\frac{\pi D^2}{4}"


For mangnin wire

Resistance=R'

Length(L')="\frac{L}{3}"

Resistivity ("\rho'" )="30\rho"

Diameter D'="\frac{D}{3}"

"R=\frac{\rho l}{A}"

Area(A')="\frac{\pi D^2}{36}"

"\frac{R'}{R}=\frac{\frac{30\rho\times{\frac{L}{3}}}{\frac{\pi D^2}{36}}}{\frac{\rho\times{L}}{\frac{\pi D^2}{4}}}"

"\frac{R'}{R}=90\\R'=90R"